OpenStudy (anonymous):

PLEASE CALCULATE DERIVATIVE WITH RESPECT TO X: SIN^-1(A/X). (Comes from Problem Set 2 Section 5A.3F) I don't understand how the chain rule generates the answer found in the solutions.

OpenStudy (amistre64):

sin-1(a/x) = y ; im assuming "a" is a constant a/x = sin(y) -a/x^2 = y' cos(y) <-- by chain rule a/(x^2 cos(y)) = y' .................................... SO what is the answer to compare with?

OpenStudy (amistre64):

sin(y) = opp/hyp = a/x a^2 + adj^2 = x^2 adj = sqrt(x^2-a^2) cos(y) = adj/hyp = sqrt(x^2 - a^2)/x \[y'=-\frac{ax} {x^2\sqrt{x^2-a^2}}\] \[y'=-\frac{a} {x\sqrt{x^2-a^2}}\] does it look anything like that?

OpenStudy (anonymous):

Thank you. Yes, that is the precisely the answer given in the solutions. I tried setting the equation to equal y and then doing implicit differentiation, but I didn't understand how to substitute for cosy- I forget that the Pythagorean theorem could be used in situations like this. Thanks also for the WA example, I think that is how the solutions derived it. Could you now explain how d/du sin^-1= 1/sqrt (1-u^2) ?

OpenStudy (amistre64):

im assuming its sin-1(u)=a this is simply definitional. the inverse sin of a ratio will tell us what angle it came from. sin-1(u) = a u/1 = sin(a) ; sin is defined as opposite/hypotenuse, so lets draw a picture. Du(u) = Du(sin(a)) 1 = cos(a) d/du d/du = 1/cos(a) ; from the drawing, cos(a) = adj/hyp: cos(a) = sqrt(1-u^2) d/du = 1/sqrt(1-u^2)

OpenStudy (amistre64):

something along those lines :)

OpenStudy (anonymous):

Thank you again, very clearly explained. I would give another medal for this question if I could. In your own experience, how often is the Pythagorean Theorem used for solving derivatives? It think I need to be more proactive in exploiting it's use.

OpenStudy (amistre64):

When dealing with the trig stuff; I find that it is always good to draw a picture so that you can visualize the process better. It comes in handy when you have to do integration.